# Embodied Image Quality Assessment for Robotic Intelligence

### v2

@@ Abstract
Image quality assessment (IQA) of user-generated content (UGC) is a critical technique for human quality of experience (QoE). However, for robot-generated content (RGC), will its image quality be consistent with the Moravec paradox and counter to human common sense? Human subjective scoring is more based on the attractiveness of the image. Embodied agent are required to interact and perceive in the environment, and finally perform specific tasks. Visual images as inputs directly influence downstream tasks. In this paper, we first propose an embodied image quality assessment (EIQA) frameworks. We establish assessment metrics for input images based on the downstream tasks of robot. In addition, we construct an Embodied Preference Database (EPD) containing 5,000 reference and distorted image annotations. The performance of mainstream IQA algorithms on EPD dataset is finally verified. The experiments demonstrate that quality assessment of embodied images is different from that of humans. We sincerely hope that the EPD can contribute to the development of embodied AI by focusing on image quality assessment. The benchmark is available at https://github.com/Jianbo-maker/EPD_benchmark.
@@
Iamge quality assessment, embodied AI, EPD
@@ I Introduction
Over the past two decades, Image Quality Assessment (IQA) has undergone significant evolution, with traditional methods aiming to mimic the Human Visual System (HVS) in image observation. However, Embodied Artificial Intelligence (EAI) faces an open challenge: understanding how image quality impacts the performance of embodied tasks. Unlike traditional assessment methods, which are limited to evaluating single, static images, EAI underscores the importance of interaction and perception within an environment. Robots perform specific tasks through interaction with their surroundings, where the image quality is crucial in determining the success of these tasks. Most approaches are offline and modular. They concentrate solely on assessing individual image quality, without considering the context of the robotic specific tasks. The robotics community has increasingly focused on the development of denoising algorithms for different types of image noise, including rain and fog removal, customized for specific environmental conditions. The development of a robot-centric evaluation methodology for embodied agents is of paramount importance for the integrated advancement of the Deep Learning (DL) technology and the robotic technology.
@@
There are inherent disparities between the human visual system and the Robot Visual System (RVS). Generally, humans prioritize the high-level semantic content of an image, whereas machines tend to analyze the low-level geometric structure. For example, in the context of a robotic task, the distortion of image luminance variation has a lesser impact on its ability to perform tasks, in contrast to the significant detriment caused by spatial distortion. Consequently, robots prioritize the consistency of image texture and structure and are less sensitive to changes in semantic information, which contrasts with human perceptions of image quality.
@@
Vision-based embodied artificial intelligence has a wide range of research prospects. However, a comprehensive framework for assessing the image quality of robotic vision systems remains undeveloped. The application of vision technology to robotics itself has inherent limitations. Vision cameras are passive sensors that passively receive information from the environment. The camera itself and its surroundings are subject to a large number of interfering factors, especially in the image collection and transmission where additional noise is introduced. Embodied intelligence designed only through ideal images is difficult to interact with the real world. The primary bottleneck hindering this research is the absence of image quality assessment for embodied AI. Considering this, this paper proposes a benchmark for quality assessment of embodied images, Fig. 1 illustrates the paradigm of the proposed benchmark. Our main contributions are as follows:
@@
Figure 2: The construction overview of the proposed EPD Benchmark. The EPD benchmark is oriented towards embodied AI and includes 25 common image distortions. There are 100 pairs of reference/distortion images for each distortion type.
  • We introduce a standardized procedure for collecting subjective labels from embodied robots. We innovatively define the subjective preferences of Embodied AI. The impact of varying image quality inputs on the robotic downstream tasks is assessed.
@@
  • We construct the embodied preference database EPD, which is the first IQA database for embodied AI. This database comprises 2,500 reference/distorted image pairs. The images are annotated via a robotic arm completing some embodied tasks. The EPD explore the internal mechanism of IQA in embodied AI.
@@ II-A Image Quality Assessment
Human-oriented image quality assessment methods can be divided into three categories, the Full-Reference (FR) IQA, the Reduced-Reference (RR) IQA, and the No-Reference (NR) IQA. The FR IQA method assumes that a distortion-free reference image exists and that the reference image information is fully available [1, 2, 3]. The RR IQA method usually assesses the quality of an image using partial information from a reference image or extracting a small number of features [4]. The NR IQA approach usually refers to the assessment of distorted image quality in the absence of the original distortion-free image [5, 6].
In addition, the benchmark for image quality assessment is a crucial issue [7, 8]. Previous work has modelled the human visual system for image quality assessment. However, there are differences in humans and embodied robots view images. To the best of our knowledge we present the first image quality assessment for Embodied Artificial Intelligence.
@@ II-B Embodied Artificial Intelligence
In contrast to traditional deep learning methods, embodied artificial intelligence emphasises the exploration of the surrounding environment by an intelligent agent, which is subject to active perception, interaction and reasoning. Existing work studies embodied intelligence from a number of different perspectives, such as embodied agent, embodied simulators, embodied perception, embodied interaction [9, 10, 11, 12, 13].
As with the development of traditional AI, embodied AI also requires a large amount of experimental data and scenarios for testing. A simulator can simulate the real physical world, subject to the limitations of experimental conditions and considerations of experimental safety. The Issac Sim [14], Gazebo [15], and PyBullet [16] provide a general-purpose simulator to support algorithm development and model training. Simulators such as AI2-THOR [17], Habitat [18], SAPIEN [19]. These platform provide a large number of scenarios that simulate real-world, task-specific scenarios, including virtual scenes and virtual objects. In this paper, the SAPIEN is used as a simulation environment to evaluate the image quality by using a robotic arm to perform specific tasks.
@@ III-A Image Collection
For distorted images, 25 common types of image distortions are set by modelling the image inputs by simulating camera noise and environmental noise. Specifically, these image distortion types are classified into 7 categories are Blus distortions, Colour distortions, Compression, Noise, Brightness change, Spatial distortions, Sharpness and contrast. A embodied task requires multiple images to work together, but the image distortion type is set to be consistent. Therefore, only the image of the initial frame is selected as the image to be evaluated. For a more extensive evaluation, 100 different initial scenes are set for each task separately. In summary, a total of 5,000 reference and distorted images are collected for the database. The construction overview of the proposed EPD benchmark is shown in Fig. 2.
@@ III-B Preference Score Collection
Based on a simulated environment, a robotic arm acts as an embodied intelligence to perform simple push and pick tasks. For the robot, different quality of image inputs have different impacts on the robot to complete the task, which is also directly related to the performance of the robot. Based on the ManiSkill platform [23], three reinforcement learning algorithms are used to complete two different tasks of push and pick boxes in the same scene. The reward value of each episodes is adopted as the performance score for evaluating the robot performance in the process of perceiving the environment and completing the embodied intelligence task. The robotic arm perceives the surrounding environment, then makes decisions based on the perception results, and finally executes the corresponding actions.
@@
Each episodes receives a reward corresponding to the score of the image, and finally the embodied Mean Opinion Score (MOS) are normalised to the range (0, 5).
IV Database Analysis
The MOS of the EPD and the subscores of the two tasks are shown in Fig. 3, and their correlation is quantified by the Spearman Rank Order Correlation Coefficient (SRCC). As can be seen from the figure, the correlation between tasks is low. Since the difficulty of both push and pick tasks is different, the reward values for the tasks are also different. The data distributions for the two tasks and the total MOS are similar, both showing an approximate normal distribution. In addition, there is a strong correlation between the overall scores and the individual tasks, with all SRCC exceeding 0.5. The EPD dataset is reliable in this category of tasks.
Figure 4: MOS score of the EPD, visualized in 25 distortion subsets. The results show the reflection of different image distortion in the perspective of the robot.
@@ IV Database Analysis
Fig. 4 illustrates the MOS distribution of the 25 distortions. From the results, it can be seen that the assessment of an image by a robot is not the same as the assessment of an image by a human. Robots are not as sensitive as humans to changes in brightness when performing tasks. The distortions in sharpness and contrast have a greater impact on robots. Essentially, embodied robots focus on the low-level pixel-level features of an image, while humans focus on the high-level semantic features of an image. Therefore, the type of distortion that affects the pixel features can seriously affect the embodied task, whereas the type of distortion that destroys the semantic integrity, which is more of a concern to the human visual system, only affects the image assessment of humans. Hence the method of image quality assessment that based on human visual system is not work well.

### v3

@@ Abstract
Image Quality Assessment (IQA) of User-Generated Content (UGC) is a critical technique for human Quality of Experience (QoE). However, does the the image quality of Robot-Generated Content (RGC) demonstrate traits consistent with the Moravec paradox, potentially conflicting with human perceptual norms? Human subjective scoring is more based on the attractiveness of the image. Embodied agent are required to interact and perceive in the environment, and finally perform specific tasks. Visual images as inputs directly influence downstream tasks. In this paper, we explore the perception mechanism of embodied robots for image quality. We propose the first Embodied Preference Database (EPD), which contains 12,500 distorted image annotations. We establish assessment metrics based on the downstream tasks of robot. In addition, there is a gap between UGC and RGC. To address this, we propose a novel Multi-scale Attention Embodied Image Quality Assessment called MA-EIQA. For the proposed EPD dataset, this is the first no-reference IQA model designed for embodied robot. Finally, the performance of mainstream IQA algorithms on EPD dataset is verified. The experiments demonstrate that quality assessment of embodied images is different from that of humans. We sincerely hope that the EPD can contribute to the development of embodied AI by focusing on image quality assessment. The benchmark is available at https://github.com/Jianbo-maker/EPD_benchmark.
@@
Image Quality Assessment, Visual Signal Processing, Image Communication, Embodied AI.
@@ I Introduction
Over the past two decades, Image Quality Assessment (IQA) has undergone significant evolution. Traditional methods are designed based on the Human Visual System (HVS) and focus on human subjective experience of images [1, 2, 3]. However, how does an Embodied Artificial Intelligence (EAI) assess image quality based on their own experience? The EAI emphasizes interaction and perception in the real environment. Robots perform specific tasks through interaction with their surroundings, where image quality is crucial in determining the success of these tasks. In addition, the robotics community is focused on the development of denoising algorithms for different types of image noise, including dehazing, deraining, and deblurring [4, 5, 6]. The assumption based on the same type of image distortion is not flexible enough, and it also lacks exploration of the relevance of image quality to embodied tasks. The development of a robot-centric image quality assessment is important for the technology of visual signal processing and embodied robots.
@@
Unlike offline machine learning, which is driven by massive amounts of internet data, embodied AI currently faces a severe bottleneck due to a lack of high-quality data. Consequently, by understanding how image quality affects downstream tasks, an embodied robot can autonomously determine which image data are beneficial for its task during acquisition and usage, and thus improving its overall performance.
@@
There are inherent disparities between the HVS and the Robot Visual System (RVS). Generally, humans prioritize the high-level semantic content of an image, whereas machines tend to analyze the low-level geometric structure. Robots prioritize the consistency of image texture and structure and are less sensitive to changes in semantic information, which contrasts with human perceptions of image quality.
Traditional IQA methods are primarily designed based on the HVS [7, 8]. These model architectures are relatively large, which are difficult to apply to embodied AI agents. Embodied robot operate under strict computational resource constraints and require real-time responses. Furthermore, the visual perception tasks of embodied AI are more focused on accurately understanding physical information within a scene, such as texture details, edge outlines, and structural integrity, rather than the aesthetic factors central to human perception.
@@
Vision-based embodied artificial intelligence has a wide range of research prospects. However, a comprehensive framework for assessing the image quality of robotic vision systems remains undeveloped. The application of vision technology to robotics itself has inherent limitations. Vision cameras are passive sensors that passively receive information from the environment. The camera itself and its surroundings are subject to a large number of interfering factors, especially in the image collection and transmission where additional noise is introduced. Embodied methods designed only through desirable image assumptions are difficult to interact effectively with the real world. The primary bottleneck that hinders this research is the absence of image quality assessment for embodied AI. Considering this, this paper proposes a benchmark and an approach for quality assessment of embodied images. Fig. 1 illustrates the paradigm of the proposed benchmark. In addition, given the inherent differences between HVS and RVS, traditional IQA approaches are not applicable to the new scenario. Therefore, we propose an embodied image quality assessment method for this new area. Our main contributions are as follows:
@@
Figure 2: The construction overview of the proposed EPD Benchmark. The EPD benchmark is oriented towards embodied AI and includes 25 common image distortions with 5 different levels of intensity. There are 100 pairs of reference/distortion images for each distortion type.
Figure 3: DMOS of the EPD, visualized in 25 distortion subsets. Different color denotes distortion strength Level 1 , Level 2, Level 3, Level 4, Level 5. Distribution results show the sensitivity of the embodied robot to each distortion varied significantly.
@@
  • We construct the embodied preference database EPD, which is the first IQA database for embodied AI. This database comprises 12,500 reference/distorted image pairs. The images are annotated via a robotic arm completing the embodied tasks without any human participation. The EPD explore the internal mechanism of IQA in embodied AI.
  • We propose the first embodied image quality assessment model, which is Multi-scale Attention Embodied Image Quality Assessment (MA-EIQA) network. MA-EIQA exceeds the performance of traditional complex IQA networks only through its lightweight structure.
@@ II-A Image Quality Assessment
Human-oriented image quality assessment methods can be divided into three categories, the Full-Reference (FR) IQA, the Reduced-Reference (RR) IQA, and the No-Reference (NR) IQA.
The FR IQA method assumes that a distortion-free reference image exists and that the reference image information is fully available. Early approaches are based primarily on image content and and structure at the pixel level. SSIM [9] combines contrast information, image brightness information, and image structural similarity information to estimate the quality of an image. A number of subsequent work has built on this foundation to continue to advance IQA, such as VW-SSIM [10], IW-SSIM [11], VIF [10], FSIM [12], GMSD [13]. In recent years, the FR-IQA approach incorporating deep learning has yielded more accurate evaluation results [14, 15].
The RR IQA method usually assesses the quality of an image using partial information from a reference image or extracting a small number of features [16]. The NR IQA approach usually refers to the assessment of distorted image quality in the absence of the original distortion-free image [17, 18]. Methods based on Natural Scene Statistics (NSS) features have been studied in the early stages. Deep learning-based algorithms for reference-free quality evaluation are gradually replacing traditional methods [19, 20].
In addition, the benchmark for image quality assessment is a crucial issue. Synthetic distortion datasets such as LIVE  [21] and KADID-10k  [22] are used to evaluate traditional image quality evaluation algorithms. In order to eliminate the gap between synthetic distorted images and distorted images in the wild, datasets such as LIVEC  [23] and KonIQ-10k  [24] have been constructed. In recent years, IQA evaluation benchmarks for multi-modal large language models (MLLMs) are proposed. Q-Bench++  [25] evaluated the performance of MLLMs on images in three aspects of perception, description, and assessment. AI-Generated Image (AGI) subjective quality assessment database and benchmark are proposed  [26, 27, 28]. R-bench  [29] evaluates the robustness of MLLMs in the real-world. Previous work has modeled the Human Visual System (HVS) for image quality assessment. However, there are differences in how humans and embodied robots view images. To the best of our knowledge we present the first image quality assessment for Embodied Artificial Intelligence.
@@ II-B Embodied Artificial Intelligence
In contrast to traditional deep learning, embodied artificial intelligence emphasizes the exploration of the surrounding environment by an intelligent agent, which is subject to active perception, interaction, and reasoning. Existing work studies embodied intelligence from a number of different perspectives, such as embodied agent, embodied simulators, embodied perception, embodied interaction [30].
With the aim of being able to actively interact with the real world beyond mere simulation in a virtual world on a computer, different robotic bodies help the AI to have a physical body that is concrete. Wheeled and tracked robots are characterized by simplicity of structure and mobility, but are limited in complex terrain [31, 32]. The biomimetic robots achieve movement in different scenarios by simulating the movement characteristics of living creatures [33, 34, 35, 36]. The mechanical structure and motion control of these robots are more complicated, and they are less reliable and stable in real scenarios. As the most mature product in industrial scenarios, fixed-base robots have gained a great deal of research due to their stability and high operational precision [37, 38]. This paper explores the extent to which image quality affects embodied tasks using a fixed-base robotic arm task for embodied robots, with more explorations of embodied intelligences to be completed subsequently.
As with the development of traditional AI, embodied artificial intelligence also requires a large amount of experimental data and scenarios for testing. In general, a simulator that can simulate the real physical world can provide an efficient and scalable experimental environment, subject to the limitations of experimental conditions and considerations of experimental safety. The Issac Sim [39], Gazebo [40], and PyBullet [41] provide a general-purpose simulator to support algorithm development and model training. Simulators such as AI2-THOR [42], Habitat [43], SAPIEN [44], etc. provide a large number of scenarios that simulate real-world, mission-specific scenarios, including virtual scenes and virtual objects. In this paper, the SAPIEN is used as a simulation environment to evaluate the image quality by using a robotic arm to perform specific tasks.
@@ III-A Image Collection
For distorted images, 25 common types of image distortions with 5 different levels of intensity are set by modelling the image inputs by simulating camera distortions and environmental distortions. Specifically, these image distortion types are classified into 7 categories are Blur distortions, Color distortions, Compression, Noise, Brightness change, Spatial distortions, Sharpness and contrast. A embodied task requires multiple images to work together, but the image distortion type is set to be consistent. Therefore, only the image of the initial frame is selected as the image to be evaluated. For a more extensive evaluation, 100 different initial scenes are set for each task separately. In summary, a total of 12,500 reference and distorted images are collected for the database. The construction overview of the proposed EPD benchmark is shown in Fig. 2.
@@ III-B Preference Score Collection
Based on a simulated environment, a robotic arm acts as an embodied intelligence to perform simple push and pick tasks. For the robot, different quality of image inputs have different impacts on the robot to complete the task, which is also directly related to the performance of the robot. Based on the ManiSkill platform [48], three reinforcement learning algorithms are used to complete two different tasks in the same scene. The reward value of each episodes is adopted as the performance score for evaluating the robot performance in the process of perceiving the environment and completing the embodied intelligence task. In addition, the evaluation of RL tasks can be based on reward values, success rates, and accuracy. However, success rates and accuracy cannot provide as detailed an assessment of task performance as reward values, which record scores for each step of a task. The robotic arm perceives the surrounding environment, then makes decisions based on the perception results, and finally executes the corresponding actions.
@@
Each episodes receives a reward corresponding to the score of the image, and finally the embodied Differential Mean Opinion Score (DMOS) are normalised to the range (0, 5).
III-C Database Analysis
@@
Fig. 3 illustrates the DMOS distribution of the 25 distortions with 5 levels. From the results, it can be seen that the assessment of an image by a robot is not the same as the assessment of an image by a human. Robots are not as sensitive as humans to changes in different levels of intensity when performing tasks. Essentially, embodied robots focus on the low-level pixel-level features of an image, while humans focus on the high-level semantic features of an image. Therefore, the type of distortion that affects the pixel features can seriously affect the embodied task. Whereas the type of distortion that destroys the semantic integrity, which is more of a concern to the human visual system, only affects the image assessment of humans. Hence the method of image quality assessment that based on human visual system is not work well in EPD.
The DMOS of the EPD and the subscores of the two tasks are shown in Fig. 4, and their correlation is quantified by the Spearman Rank Order Correlation Coefficient (SRCC). As can be seen from the figure, the correlation between tasks is low. Since the difficulty of both push and pick tasks is different, the reward values for the tasks are also different. The data distributions for the two tasks and the total DMOS are similar, both showing an approximate normal distribution. In addition, there is a strong correlation between the overall scores and the individual tasks, with all SRCC exceeding 0.5. The EPD dataset is reliable in this category of tasks.
[…]
With the intention of exploring the differences between embodied intelligence and humans with respect to IQA, experiments compared the correlation between human subjective scores and scores on embodied tasks. The correlation is quantified by the Pearson Linear Correlation Coefficient (PLCC). A human subjective experiment is conducted to obtain the MOS of the EPD database. Images obtained from the embodied task were evaluated by 15 experienced experts to obtain a human MOS. Fig. 5 illustrates the correlation between humans and robots on the EPD database. The correlation between human and embodied intelligence preference scores for images is low, with a correlation coefficient of 0.2116, 0.1778, 0.1297 respectively. This also confirms that there is a gap between machines and humans regarding image quality, and therefore the traditional HVS-based IQA method not be adapted to embodied intelligence.
IV Proposed Methed
The task of image quality assessment for embodied AI face specific challenges in terms of real-time model perception. On one hand, the computational resources of hardware platforms on which embodied robots are limited, requiring IQA models to be lightweight and efficient. On the other hand, the visual systems of embodied AI are more concerned with physical world information that directly impacts their decision and execution. For embodied robots, the object texture details, scene structure, and potential motion blur are more important, rather than the aesthetic quality preferred by the human visual system.
Towards these requirements, this paper proposes a novel multi-scale attention embodied image quality assessment, called MA-EIQA. This network utilizes an efficient multi-scale feature encoder to capture rich features crucial for robotic tasks. It also introduces an embodied attention module to dynamically focus on the most informative feature regions. Finally, a linear layer is used to regress an accurate quality score.
IV-A Multi-scale Attention Embodied Image Quality Assessment
The overall architecture of the proposed MA-EIQA network is shown in Fig. 6 and consists of two main components: a multi-scale feature encoding module and an embodied attention module.
The specific workflow is as follows: First, an input image $I$ is fed into a pre-trained ResNet50 [49] backbone to extract foundational, multi-level visual features. Subsequently, the feature maps from different stages of the ResNet50 are passed to the multi-scale feature encoder. Inspired by the design of the Path Aggregation Network (PANet) [50], we generates an information-rich, aggregated feature map $F\_{E}$ through bi-directional path fusion.
Next, $F\_{E}$ is sent to the embodied attention module, which refines the features through a series of channel and spatial attention mechanisms to produce an attention-weighted feature map $F\_{A}$.
Finally, the feature map $F\_{A}$ is flattened and passed through two consecutive fully connected linear layers to regress the final image quality score $y$.
The loss function is used to measure the difference between the model predicted score $y$ and the DMOS, denoted as $\hat{y}$, which is annotated by an embodied ”expert”. Given its penalty for larger errors and its smoothness in gradients, we select the Mean Squared Error (MSE Loss) as the objective function for model optimization. Its definition is as follows:
$L_{MSE}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2},$ (4)
where $N$ is the number of samples in a batch, $y_{i}$ is the predicted score for the $i$-th image, and $\hat{y}_{i}$ is the corresponding ground-truth DMOS.
Figure 6: Framework of the proposed method. MA-EIQA is the first embodied No-Reference IQA model, which consists of two components: multi-scale feature extraction and embodied attention module. The input image II goes through feature extraction, and then $F_{E}$ is fed into the attention module. Finally, the extracted features $F_{A}$ are subjected to a linear layer to obtain the final prediction score.
IV-B Multi-Scale Feature Encoder
When interacting with the physical world, embodied AI agents need to understand both the macro-level layout (semantic information) and micro-level details (texture, edge information) of a scene. In embodied manipulation tasks, it is crucial to perceive both the semantic attributes of the task and learn useful information from textural features. A single-scale feature map struggles to satisfy both of these needs simultaneously.
To address this issue, we propose an efficient multi-scale feature encoder. Aiming to achieve deep, cross-scale feature fusion with a lightweight manner. The central idea is to aggregate features from different levels through both a bottom-up and a top-down path. This helps to effectively fuse high-level semantic information with low-level precise localization details. The ResNet50 [49] is selected as the backbone because it can provide rich multi-scale features.
The input features are from different stages of the ResNet50, $C_{2},C_{3},C_{4},C_{5}$}, where $C_{i}$ represents the output of the $i$-th stage. The subsequent process involves two information propagation paths:
Top-Down Feature Pyramid Path. This path is responsible for transmitting high-level semantic abstractions to lower-level layers. The hierarchical feature pyramid construction initiates from the deepest feature map ($C_{5}$), an initial feature pyramid $P_{2},P_{3},P_{4},P_{5}$} is generated by progressively upsampling and using 1x1 convolutions (for channel matching) to fuse with lower-level features. This process can be expressed as:
$P_{i}=f_{up}(P_{i+1})+T_{i}(C_{i}),\quad i\in 2,3,4,$ (5)
where $f_{up}(\cdot)$ denotes an upsampling operation (such as bilinear interpolation), and $\mathcal{T}_{i}(\cdot)$ represents a 1x1 convolutional layer used for adjusting channel dimensions and feature projection.
Bottom-Up Feature Enhancement Path. To compensate for the potential dilution of low-level feature information that can occur during long-range propagation in the top-down path, an additional bottom-up path is constructed. Starting from the lower-level pyramid feature map ($P_{2}$), this path utilizes the 3×3 convolutions (stride=2) for hierarchical downsampling. This path progressively enhance higher-level feature maps with rich spatial details from lower layers, producing the final feature pyramid $N_{2},N_{3},N_{4},N_{5}$}:
$N_{i}=f_{down}(N_{i-1})+P_{i},\quad i\in 3,4,$ (6)
where $N_{2}=P_{2}$, and $f_{down}(\cdot)$ represents a 3x3 convolutional layer with a stride of 2, used for downsampling and feature transformation.
The final output feature F_EF\_E from this module is able to contain both rich textural and edge details as well as high-level semantic information, providing a high-quality input for the subsequent attention module.
IV-C Embodied Attention Module
Not all features are equally important for the final quality assessment. To enable the model to adaptively focus on these critical features that are useful to robots, we propose an Embodied Attention Module. The structure of this module is based on the Convolutional Block Attention Module (CBAM) [51]. The module takes the feature $F_{E}\in\mathbb{R}^{C\times H\times W}$ from the previous stage as input and outputs an attention-weighted feature map $F_{A}$. This module refines and weights features from both channel and spatial dimensions through a series of two sub-modules: Channel Attention and Spatial Attention.
Channel Attention Module. This module aims to model the importance of different feature channels. It aggregates the spatial information of each channel using both global average pooling (AvgPool) and global maximum pooling (MaxPool). Both are then fed into a shared, lightweight multi-layer perceptron (MLP), and the final channel attention weights $M_{c}\in\mathbb{R}^{C\times 1\times 1}$ are generated through a Sigmoid activation function.
$M_{c}(F_{E})=\sigma(\mathrm{MLP(AvgPool}(F_{E}))+\mathrm{MLP(MaxPool}(F_{E}))),$ (7)
where $\sigma$ denotes the Sigmoid function. The resulting channel attention weights are then broadcast and multiplied element-wise with the original feature map to obtain the channel-weighted feature map $F^{\prime}$:
$F^{\prime}=M_{c}(F_{E})\otimes F_{E}$ (8)
Spatial Attention Module. This module follows the channel attention and is designed to identify important regions within the feature map. It first applies both average pooling and max pooling along the channel dimension of the channel-weighted feature map, $F^{\prime}$, to generate two 2D spatial descriptors. These two descriptors are then concatenated and fused via a standard convolutional layer (i.e., a 7x7 convolution). Finally, a 2D spatial attention map, $M_{s}\in\mathbb{R}^{1\times H\times W}$, is generated using a Sigmoid activation function.
$M_{s}(F^{\prime})=\sigma(f^{7\times 7}([\mathrm{AvgPool}*ch(F^{\prime});\mathrm{MaxPool}*ch(F^{\prime})])),$ (9)
where $f^{7\times 7}$ denotes a convolution operation with a $7\times 7$ kernel, and $\cdot$$\cdot$] represents a concatenation operation. Similarly, the resulting spatial attention map is broadcasted and multiplied element-wise with $F^{\prime}$ to obtain the final refined feature map, $F_{A}$:
$F_{A}=M_{s}(F^{\prime})\otimes F^{\prime}$ (10)
By processing features through the embodied attention module in series, the model is able to suppress irrelevant background and noise, while simultaneously enhancing the feature representations most relevant to image quality. This provides a more focused feature representation for the subsequent quality score regression.